OpenStudy (anonymous):

Given g(x) = x2 + 5 and h(x) = x − 3, evaluate the composite function. h[g(x)]

OpenStudy (anonymous):

g(x) = x^2 + 5 h(g(x) =h (x^2 + 5) = x^2 + 5 -3 = x^2 + 2

hero (hero):

cassyynn, I posted a solution to your first question

OpenStudy (anonymous):

thank you

hero (hero):

the one where you had to find (f*g)(1)

OpenStudy (anonymous):

sweet thank you